OpenStudy (anonymous):

What is the sum of a 22-term arithmetic sequence where the first term is 57 and the last term is -27

OpenStudy (anonymous):

you need to know the numbers between

OpenStudy (anonymous):

no just the sum of the numbers

OpenStudy (anonymous):

ight then ig add all the numbers

OpenStudy (anonymous):

so wat are the numbers in between..

OpenStudy (anonymous):

well. I think since there are 22 terms and the first one is 57 and the last one is -27 it means that: \[\LARGE a_1=57 \quad \quad ,\quad \quad a_{22}=-27\] \[\LARGE a_{22}=a_1+(22-1)d\] \[\LARGE -27=57+21d\] ... whoaa @eliassaab O_O :)

OpenStudy (anonymous):

so wat would be my final answer my choices are A) 240 B) 270 C) 300 D) 330

OpenStudy (anonymous):

@Kreshnik, Solve for d, d=-4

OpenStudy (anonymous):

...so it means that d=-4 \[\LARGE Sn=\frac n2 [ 2\cdot a_1+(n-1)d]\] \[\LARGE S_{22}=\frac{22}{2} [ 2\cdot 57+(22-1)(-4)]\] \[\LARGE S_{22}=11 [ 114-84]\] \[\LARGE S_{22}=11 [30 ]\] \[\LARGE S_{22}=330\]

OpenStudy (anonymous):

Here is the sequence {57, 53, 49, 45, 41, 37, 33, 29, 25, 21, 17, 13, 9, 5, 1, -3, -7, \ -11, -15, -19, -23, -27} And the sum is 330

OpenStudy (anonymous):

@eliassaab well done. Great work ! ;)

OpenStudy (anonymous):

thanks guys ;)